OpenStudy (anonymous):

Evaluate. x^2+11x+30/x^2+12x+36 / x^2+8x+15/x^2+3x-18

OpenStudy (campbell_st):

well dividing by a fraction invert and multiply so the problem is \[\frac{x^2 + 11x + 30}{x^2 + 12x + 36} \times \frac{x^2 + 8x + 15}{x^2 + 3x - 18}\] now factorise the quadratics and eliminate common factors \[\frac{(x+5)(x+6)}{(x+6)^2} \times \frac{(x + 5)(x+3)}{(x+6)(x-3)}\] so (x+6) and (x-3) are common factors leaving \[\frac{(x+5)^2}{(x+6)^2} =\frac{x^2 + 10x + 25}{x^2 + 12x+36}\]

OpenStudy (anonymous):

thatt is a division sign

OpenStudy (campbell_st):

where... between the fractions...?

OpenStudy (campbell_st):

rule for fractions divide by a fraction... find the reciprocal and multiply

OpenStudy (anonymous):

my choices are

OpenStudy (anonymous):

(x+5)(x+3)/(x-3)

OpenStudy (anonymous):

x-3/x+3

OpenStudy (anonymous):

x+6/x+5

OpenStudy (campbell_st):

well you can't evaluate the problem as its not equal to anything... and you don't have any numbers based on your question... my solution seems sound